Hi everyone, I have two problems to solve:
x^2+3x-4=0
x^2+2x-8=0
Thank you,
Gina

Answer by nerdybill(7384)

Both can be solved by factoring.
x^2+3x-4=0
(x+4)(x-1) = 0
x = {-4, 1}
.
x^2+2x-8=0
(x+4)(x-2) = 0
x = {-4, 2}